OpenStudy (anonymous):

3/4=9/x I need to know how to solve this not just the answer please.

OpenStudy (anonymous):

Cross-multiply and divide by 3

OpenStudy (phi):

you can cross multiply?

OpenStudy (cwrw238):

find lcd then multiply through b this to eliminate the fractions then solve f or x

OpenStudy (anonymous):

I did the cross multiplying. 3*x=3x & 4*9=36 That gave me 3x=36 Then i did 3x/3= x & 36/3= 12 so i got x=12 Is that right?

OpenStudy (cwrw238):

dead right

OpenStudy (anonymous):

YAY!! THX!!!!

OpenStudy (cwrw238):

cross multiplying works when you have 1 or 2 fractions - if there are more than 2 fractions the method i mentioned can be used